Calculate the number the number of moles of solute contained in 3.0 L of 4.0 M Solution
1
Expert's answer
2017-11-06T11:26:06-0500
n(solute) = c(solute) × V(solution) = 4.0 × 3.0 = 12.0 mol

Answer:
12.0 mol.